# Bigger is not Always Better: Scaling Properties of Latent Diffusion Models

### v1

@@ 3.4 Scaling sampling-efficiency in different samplers
To assess the generalizability of observed scaling trends in sampling efficiency, we compared scaled LDM performance using different diffusion samplers. In addition to the default DDIM sampler, we employed two representative alternatives: the stochastic DDPM sampler [16] and the high-order DPM-Solver++ [35].
@@ 3.5 Scaling downstream sampling-efficiency
Refer to captionRefer to caption
Figure 13: Distillation improves text-to-image performance and scalability. Left: Distilled Latent Diffusion Models (LDMs) consistently exhibit lower (better) FID scores compared to their undistilled counterparts across varying model sizes. The consistent acceleration factor (approx. 5×5$\times$5 ×) indicates that the benefits of distillation scale well with model size. Right: Distilled models using only 4 sampling steps achieve FID scores comparable to undistilled models using significantly more steps. Interestingly, at a sampling cost of 7, the distilled 866M model performs similarly to the smaller, undistilled 83M model, suggesting improved efficiency. See the supplement for visual examples demonstrating the quality improvements achieved through distillation.
@@
We have featured the scaling sampling-efficiency of latent diffusion models, which demonstrates that smaller model sizes exhibit higher sampling efficiency. A notable caveat, however, is that smaller models typically imply reduced modeling capability. This poses a challenge for recent diffusion distillation methods [36, 59, 65, 61, 14, 41, 37, 29] that heavily depend on modeling capability. One might expect a contradictory conclusion and believe the distilled large models sample faster than distilled small models. In order to demonstrate the sampling efficiency of scaled models after distillation, we distill our previously scaled models with conditional consistency distillation [65, 41] on text-to-image data and compare those distilled models on their optimal performance. Please see our supplement for more distillation details.
To elaborate, we test all distilled models with the same 4-step sampling, which is shown to be able to achieve the best sampling performance; we then compare each distilled model with the undistilled one on the normalized sampling cost. We follow the same practice discussed in Section 3.3.1 for selecting the optimal CFG rate and compare them under the same relative inference cost. The results shown in the left part of Fig. 13 demonstrate that distillation significantly improves the generative performance for all models in 4-step sampling, with FID improvements across the board. By comparing these distilled models with the undistilled models in the right part of Fig. 13, we demonstrate that distilled models outperform undistilled models at the same sampling cost. However, at the specific sampling cost, i.e., sampling cost $\approx\texttt{8}$, the smaller undistilled 83M model still achieves similar performance to the larger distilled 866M model. The observation further supports our proposed scaling sampling-efficiency of LDMs, which still holds under the circumstance of diffusion distillation.
@@ 4 Conclusion
In this paper, we investigated scaling properties of Latent Diffusion Models (LDMs), specifically through scaling model size from 39 million to 5 billion parameters. We trained these scaled models from scratch on a large-scale text-to-image dataset and then finetuned the pretrained models for downstream tasks. Our findings unveil that, under identical sampling costs, smaller models frequently outperform larger models, suggesting a promising direction for accelerating LDMs in terms of model size. We further show that the sampling efficiency is consistent in multiple axes. For example, it is invariant to various diffusion samplers (stochastic and deterministic), and also holds true for distilled models. We believe this analysis of scaling sampling efficiency would be instrumental in guiding future developments of LDMs, specifically for balancing model size against performance and efficiency in a broad spectrum of practical applications.
Limitations.
This work utilizes visual quality inspection alongside established metrics like FID and CLIP scores. We opted to avoid human evaluations due to the immense number of different combinations needed for the more than 1000 variants considered in this study. However, it is important to acknowledge the potential discrepancy between visual quality and quantitative metrics, which is actively discussed in recent works [73, 19, 7]. Furthermore, claims regarding the scalability of latent models are made specifically for the particular model family studied in this work. Extending this analysis to other model families, particularly those incorporating transformer-based backbones [47, 43, 38], would be a valuable direction for future research.
@@ 5 Acknowledgments
We are grateful to Keren Ye, Jason Baldridge for their valuable feedback. We also extend our gratitude to Shlomi Fruchter, Kevin Murphy, Mohammad Babaeizadeh, and Han Zhang for their instrumental contributions in facilitating the initial implementation of the latent diffusion models.
Appendix 0.A Scaling the text-to-image performance
@@
Appendix 0.B Scaling downstream performance
@@
Appendix 0.C Scaling sampling-efficiency in distilled LDMs
Diffusion distillation methods for accelerating sampling are generally derived from Progressive Distillation (PD) [59] and Consistency Models (CM) [65]. In the main paper, we have shown that CoDi [41] based on CM is scalable to different model sizes. Here we show other investigated methods, i.e., guided distillation [44], has inconsistent acceleration effects across different model sizes. Fig. 19 shows guided distillation results for the 83M and 223M models respectively, where s16 and s8 denote different distillation stages. It is easy to see that the performance improvement of these two models is inconsistent.
@@ Appendix 0.C Scaling sampling-efficiency in distilled LDMs
Appendix 0.D Scaling the sampling-efficiency

### v2

@@ 3.4 Scaling sampling-efficiency in different samplers
To assess the generalizability of observed scaling trends in sampling efficiency, we compared scaled LDM performance using different diffusion samplers. In addition to the default DDIM sampler, we employed two representative alternatives: the stochastic DDPM sampler (Ho et al., 2020) and the high-order DPM-Solver++ (Lu et al., 2022b).
@@ 3.5 Scaling downstream sampling-efficiency
Figure 13: Distillation improves text-to-image performance and scalability. Left: Distilled Latent Diffusion Models (LDMs) consistently exhibit lower (better) FID scores compared to their undistilled counterparts across varying model sizes. The consistent acceleration factor (approx. 5×5$\times$5 ×) indicates that the benefits of distillation scale well with model size. Right: Distilled models using only 4 sampling steps achieve FID scores comparable to undistilled models using significantly more steps. Interestingly, at a sampling cost of 7, the distilled 866M model performs similarly to the smaller, undistilled 83M model, suggesting improved efficiency.
@@
We have featured the scaling sampling-efficiency of latent diffusion models, which demonstrates that smaller model sizes exhibit higher sampling efficiency. A notable caveat, however, is that smaller models typically imply reduced modeling capability. This poses a challenge for recent diffusion distillation methods (Luhman & Luhman, 2021; Salimans & Ho, 2022; Song et al., 2023; Sauer et al., 2023b; Gu et al., 2023; Mei et al., 2024a; Luo et al., 2023; Lin et al., 2024) that heavily depend on modeling capability. One might expect a contradictory conclusion and believe the distilled large models sample faster than distilled small models. In order to demonstrate the sampling efficiency of scaled models after distillation, we distill our previously scaled models with conditional consistency distillation (Song et al., 2023; Mei et al., 2024a) on text-to-image data and compare those distilled models on their optimal performance.
To elaborate, we test all distilled models with the same 4-step sampling, which is shown to be able to achieve the best sampling performance; we then compare each distilled model with the undistilled one on the normalized sampling cost. We follow the same practice discussed in Section 3.3.1 for selecting the optimal CFG rate and compare them under the same relative inference cost. The results shown in the left part of Fig. 13 demonstrate that distillation significantly improves the generative performance for all models in 4-step sampling, with FID improvements across the board. By comparing these distilled models with the undistilled models in the right part of Fig. 13, we demonstrate that distilled models outperform undistilled models at the same sampling cost. However, at the specific sampling cost, i.e., sampling cost $\approx\texttt{8}$, the smaller undistilled 83M model still achieves similar performance to the larger distilled 866M model. The observation further supports our proposed scaling sampling-efficiency after diffusion distillation.
@@ 4 Conclusion
In this paper, we investigated scaling properties of Latent Diffusion Models (LDMs), specifically through scaling model size from 39 million to 5 billion parameters. We trained these scaled models from scratch on a web-scale text-to-image dataset and then finetuned the pretrained models for downstream tasks. Our findings unveil that, under identical sampling costs, smaller models frequently outperform larger models, suggesting a promising direction for accelerating LDMs in terms of model size. We further show that the sampling efficiency is consistent in multiple axes. For example, it is invariant to various diffusion samplers (stochastic and deterministic), and also holds true for distilled models. We believe this analysis of scaling sampling efficiency would be instrumental in guiding future developments of LDMs, specifically for balancing model size against performance and efficiency in a broad spectrum of practical applications.
Limitations and future work.
This work utilizes visual quality inspection alongside established metrics like FID and CLIP scores. We opted to avoid human evaluations due to the immense number of different combinations needed for the more than 1000 variants considered in this study. However, it is important to acknowledge the potential discrepancy between visual quality and quantitative metrics, which is actively discussed in recent works (Zhang et al., 2021; Jayasumana et al., 2024; Cho et al., 2023).
Claims regarding the scalability of latent diffusion models are made specifically for the particular model family studied in this work (Rombach et al., 2022). Extending this analysis to other model families, particularly those incorporating transformer-based backbones such as DiT (Peebles & Xie, 2023; Mei et al., 2023), SiT (Ma et al., 2024), MM-DiT (Esser et al., 2024), and DiS (Fei et al., 2024), and cascaded diffusion models such as Imagen3 (Baldridge et al., 2024) and Stable Cascade (Pernias et al., 2023), would be a valuable direction for future research.
@@ 5 Acknowledgments
Vishal M. Patel was supported by NSF CAREER award 2045489. We are grateful to Keren Ye, Jason Baldridge, Kelvin Chan for their valuable feedback. We also extend our gratitude to Shlomi Fruchter, Kevin Murphy, Mohammad Babaeizadeh, and Han Zhang for their instrumental contributions in facilitating the initial implementation of the latent diffusion models.
Appendix A Scaling the text-to-image performance
@@
Appendix B Scaling downstream performance
@@
Appendix C Scaling sampling-efficiency in distilled LDMs
Diffusion distillation methods for accelerating sampling are generally derived from Progressive Distillation (PD) Salimans & Ho (2022) and Consistency Models (CM) Song et al. (2023). In the main paper, we have shown that CoDi Mei et al. (2024a) based on CM is scalable to different model sizes. Here we show other investigated methods, i.e., guided distillation Meng et al. (2023), has inconsistent acceleration effects across different model sizes. Fig. 19 shows guided distillation results for the 83M and 223M models respectively, where s16 and s8 denote different distillation stages. It is easy to see that the performance improvement of these two models is inconsistent.
@@ Appendix C Scaling sampling-efficiency in distilled LDMs
Appendix D Scaling the sampling-efficiency
@@
Appendix E Scaling interpretability of text prompt interpolatation
Text prompt interpolation is widely recognized as a way to evaluate the interpretability of text-to-image models in recent works (Li et al., 2024; Park et al., 2023). In Figure 24, we show the text-prompt interopolation results of models in different sizes and visualize their sampling results. Specifically, we use two distinct prompts $A$ and $B$ and interpolate their CLIP embeddings as $\alpha A+(1-\alpha)B,\alpha\in[0,1]$, to generate intermediate text-to-image results. A clear pattern emerges: larger models leads to more semantically coherent and visually plausible interpolations compared to their smaller counterparts. The figure demonstrates the 2B model’s superior ability to accurately interpret interpolated prompts, as evidenced by its generation of a tablet computer with a touch pen.
Refer to caption 145M $\alpha=1.0$
Refer to caption 145M $\alpha=0.8$
Refer to caption 145M $\alpha=0.5$
Refer to caption 145M $\alpha=0.2$
Refer to caption 145M $\alpha=0.0$
Refer to caption 318M $\alpha=1.0$
Refer to caption 318M $\alpha=0.8$
Refer to caption 318M $\alpha=0.5$
Refer to caption 318M $\alpha=0.2$
Refer to caption 318M $\alpha=0.0$
Refer to caption 430M $\alpha=1.0$
Refer to caption 430M $\alpha=0.8$
Refer to caption 430M $\alpha=0.5$
Refer to caption 430M $\alpha=0.2$
Refer to caption 430M $\alpha=0.0$
Refer to caption 866M $\alpha=1.0$
Refer to caption 866M $\alpha=0.8$
Refer to caption 866M $\alpha=0.5$
Refer to caption 866M $\alpha=0.2$
Refer to caption 866M $\alpha=0.0$
Refer to caption 2B $\alpha=1.0$
Refer to caption 2B $\alpha=0.8$
Refer to caption 2B $\alpha=0.5$
Refer to caption 2B $\alpha=0.2$
Refer to caption 2B $\alpha=0.0$
(a) Prompt $A$: “a old computer”. Prompt $B$: “a fancy mobile phone”.
Figure 24: We visualize the text-prompt interpolation results of scaled models in different sizes. Each row shows the results of the same model with different interpolation fraction $\alpha A+(1-\alpha)B$. All results are sampled with the same 20-step DDIM sampler and CFG of 7.5.